Ticket sales. Tickets for a concert were sold to adults for
$3 and to students for $2. If the total receipts were $824
and twice as many adult tickets as student tickets were
sold, then how many of each were sold?

Answer by checkley77(12844)

A=2S
3A+2S=824
3*2S+2S=824
6S+2S=824
8S=824
S=824/8
S=103 STUDENT TICKETS WERE SOLD.
2*103=206 ADULT TICKETS WERE SOLD.
PROOF:
3*206+2*103=824
618+206=824
824=824